OpenStudy (anonymous):

What is the distance between the origin and the point (8, -17)? If necessary, round your answer to two decimal places

OpenStudy (anonymous):

\[\sqrt{(8-0)^{2} +(-17-0)^{2}} \]

OpenStudy (mertsj):

yep.

terenzreignz (terenzreignz):

I don't get it, what's wrong? :D Everything seems to be in order...

OpenStudy (anonymous):

ok is my answer going to be a decimal cuz i got 18.79 i feel its wrong

terenzreignz (terenzreignz):

Well, no, it isn't :D

terenzreignz (terenzreignz):

I mean, it isn't wrong, that is...

OpenStudy (mertsj):

\[\sqrt{64+289}=\sqrt{353}=18.8\]

OpenStudy (anonymous):

ok i was right :D

OpenStudy (mertsj):

yep